A rocket engine has a mass (fuel) of 4.1 Kg. A thrust of 10 Newtons in 0.2 sec with a total impulse of 2.5 Newton – sec. Find the specific impulse.
Answer #1

Total impulse

ITotal = mV

=>V=2.5/4.1 =0.61 m/s

Specific Impulse is given by

iSP =V/g =0.61/9.81 =0.06216 Newton-second